### v3

@@ Discussion
I.E.H., S.E., and B.M. designed the study. I.E.H. and S.E. were responsible for data collection, data analysis, model construction, model validation, figure preparation, and manuscript writing. F.A., A.G.S., S.N.E., and I.D. managed data anonymization and annotation. M.F.D. contributed to model construction. C.W., W.D., and K.B. contributed to the external evaluation. O.F.D. contributed to the graphical user interface and manuscript writing. M.X. and T.A. contributed to the creation of anatomical segmentation labels. B.W., E.S., M.S., E.B.E., A.A., A.S., B.L., C.B., and M.K.O. provided knowledge support. B.M. supervised the all study. All authors reviewed the manuscript and approved the final version.
Acknowledgments
We extend our gratitude to the Helmut Horten Foundation for their invaluable support, which made this work possible. Additionally, we would like to express our sincere appreciation to Istanbul Medipol University Mega Hospital for their support and provision of data. C.B. additionally receives research support from the Promedica Foundation, Chur, Switzerland. C.W., W.D., and K.B. receive funding from R01HL141813 (An Integrative Radiogenomic Approach to Design Genetically-Informed Image Biomarkers for Characterizing COPD), the Hariri Institute for Computing’s Junior Faculty Fellows Program, and the Hariri Institute Focused Research Program. Figures 1, 3, 5, and 6 are created with BioRender.com.

### v4

@@ Acknowledgments
I.E.H., S.E., and B.M. designed the study. I.E.H. and S.E. handled data collection, data analysis, model construction and validation, figure preparation, and manuscript writing. F.A., A.G.S., S.N.E., and I.D. managed data anonymization and annotation. M.F.D. contributed to model construction. C.W., W.D., and K.B. carried out the external evaluation. O.F.D. developed the graphical user interface and assisted with manuscript writing. M.X. and T.A. generated anatomical-segmentation labels. S.S. maintained the model repository. B.H. assisted in preparing NIfTI files with metadata and integrating nodule-segmentation masks into the report-generation pipeline. M.P. assessed report-generation performance. B.K. provided technical expertise and contributed to the manuscript writing. H.D., B.W., E.S., M.S., E.B.E., A.A., A.S., A.K., Z.L., B.L., C.B., and M.K.O. provided domain-knowledge support. B.M. supervised the entire study. All authors reviewed and approved the final manuscript.
Competing interests
The authors declare no competing interests.